# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 9.2.2 FirstDegreeTrendSB1 and SecondDegreeTrendSB1 classes
Fig. 33 illustrates the distribution of the first derivative normalised by its uncertainty. The distribution is very similar for the two classes and presents an almost perfect symmetry, as expected. The deficit of solutions around the zero derivative is due to the fact that these solutions are not included because they correspond to constant objects. A lot of the solutions are well above . A few well-defined derivatives are further away up to .
@@ 9.3 Validation against CU7 photometric periods
Some intrinsic variables can exhibit RV variations that could be mistaken as the effect of an orbital motion. As explained in Sect. 10.1.1, a sample of securely identified Cepheids (333, Ripepi et al. 2023) and RR Lyrae (147, Clementini et al. 2023) have also an SB1 solution. They should be eventually excluded from our SB-subcatalogue but they also offer the unique opportunity to have an independent test of the spectroscopic periods by comparing them to the photometric ones. Indeed, we compared our spectroscopic periods to those determined from DR3 photometric data for the 147 RR Lyrae and the 333 Cepheids. The difference is expressed as [P(spectroscopy)–P(photometry)]/, where is the square root of the sum in quadrature of both uncertainties. We find a recovery success rate to within 3 of 95% for the RR Lyrae’s. It is not only lower for the Cepheids (85%), but the outliers are also much more severely discrepant (Fig. 35). We find that the solution efficiency is better for the RR Lyrae’s than for the Cepheids.
@@ 9.4.1 Comparison against Griffin’s results
SB1: 83 systems
SB1C: 0 system
TrendSB1: 10 systems
SB2: 1 system
SB2C: 0 system
@@ 9.4.2 Comparison against SB9 catalogue
As a next step, the analysis above was repeated for the whole SB9. Although the sample is now considerably larger, it is much less homogeneous and contains solutions of generally lower quality. The most recent orbit was selected when a given system had multiple entries in the catalogue. As described above, visual/multiple systems were also excluded. Six SB2 orbits were replaced by those determined by Lester (2020). This validation catalogue contains the orbits of 1750 SB1’s and 647 SB2’s.
The breakdown of DR3 spectroscopic solution types is as follows for SB1’s in the SB9 catalogue:
SB1: 226 systems
SB1C: 0 system
TrendSB1: 45 systems
SB2: 3 systems
SB2C: 0 system
@@
All but one of the 45 systems with a trend solution have a very long period. Eighteen and two systems among the SB2 validation catalogue have received a SB1 and TrendSB1 solution, respectively.
@@ 9.4.3 Comparison against APOGEE
Orbital parameters for 20,000 SB1’s have been released as part of APO Galactic Evolution Experiment (APOGEE)131313http://www.sdss3.org/surveys/apogee.php DR16 (Price-Whelan et al. 2020a, hereafter PW20). Our validation sample is their ’Gold Sample’ that contains 1032 systems with high-quality, unimodal posterior samplings.
The breakdown of DR3 spectroscopic solution types is as follows for the SB1’s in PW20:
SB1: 126 systems
SB1C: 0 system
TrendSB1: 4 systems
SB2: 3 systems
SB2C: 0 system
@@ 9.4.4 Comparison against WISE
A total of 56,000 short-period ( 10 d) binary systems observed by the Wide-field Infrared Survey Explorer (WISE)141414https://www.jpl.nasa.gov/missions/wide-field-infrared-survey-explorer-wise have been discussed by Petrosky et al. (2021). Most of them are (near-)contact EBs. However, as will be shown below, a non-negligible amount are classical pulsators, especially RR Lyrae stars. The infrared photometric measurements are made through the W1 filter centred at about 3.4 m. A subset of 49,465 systems can be unambiguously associated to a DR3 source151515http://cdn.gea.esac.esa.int/Gaia/gedr3/cross_match/. This sample allows us to test the reliability of our orbital periods for very close binaries, although other orbital parameters are not available.
The breakdown of DR3 spectroscopic solution types is as follows for the SB1’s in Petrosky et al. (2021):
SB1: 704 systems
SB1C: 3 systems
TrendSB1: 12 systems
SB2: 69 systems
SB2C: 23 systems
Without any DR3 spectroscopic deterministic solutions: 48,654 systems.
@@ 9.4.5 Comparison against ASAS-SN
The physical parameters of 35,464 detached eclipsing binaries (DEBs) observed in the framework of the All-Sky Automated Survey for Supernovae (ASAS-SN; Kochanek et al. 2017)161616https://www.astronomy.ohio-state.edu/asassn/ have been determined by Rowan et al. (2022). A total of 35,307 systems have a Gaia DR3 ID. The data are based on the analysis of - and -band light curves. The orbital periods nicely bridge the gap between the WISE sample of close binaries and other more general catalogues, such as APOGEE or the SB9 (Fig. 36). Contrary to WISE, other orbital elements (e.g. eccentricity) are provided besides the period, although without any uncertainties. For , we assume a typical relative uncertainty of 10-6 (e.g. Holanda et al. 2018).
@@
SB1: 626 systems
SB1C: 3 systems
TrendSB1: 6 systems
SB2: 92 systems
SB2C: 35 systems
Without any DR3 spectroscopic deterministic solutions: 34,545 systems.
@@ 9.4.6 Summary
As discussed above, the success rate for the main orbital parameters (, , , and ) is usually well above 80% for the whole set of validation catalogues considered here (Griffin, SB9, APOGEE, WISE, and ASAS-SN). As expected, the performance increases for relatively short-period systems for which at least one orbital period is covered by the observations (about 2.5 years for the RVS spectroscopic processing data set considered here). In this case, the recovery rate commonly exceeds 85-90%. The WISE sample of (near-)contact binaries allows us to investigate the performance of our period search at high frequencies. We recover the WISE photometric periods at a level of about 80-85%. It increases to about 90% for the DEBs with wider orbits discussed by Rowan et al. (2022). One can also note the excellent success rate for the pulsation periods of the RR Lyrae stars that are below one day (Sect. 9.3). The good agreement on the period with both EB data sets advocates for a deeper study using these data. Their combination will certainly result in an extensive analysis that could reveal various kinds of objects such as EW binaries and/or ellipsoidal variables. Such a task is beyond the scope of the present paper. Our analysis globally suggests that the orbital periods below about 10 days are generally robust provided the solution significance is not too low. We find a roughly constant success rate at the level of about 90-95% once the significance exceeds 15-20. However, although they can often be regarded as compatible with zero within the uncertainties, it appears that the eccentricities may be overestimated even if the significance is larger.
@@
In this respect, the fact that only a relatively small fraction of the SB1’s in the validation catalogues have a spectroscopic deterministic DR3 solution is the consequence of the various filtering steps, both at CU4 and CU6 levels, described previously. In addition, we note that for the SB9 and Griffin’s sample, only a subset of the reference catalogue was used. The case of the SB9, and the steps that led to the relatively small overlap with respect to DR3 (see Sect. 9.4.2) are discussed in detail in Appendix G. Similar arguments apply to the other external catalogues and explain the limited number of benchmark binaries with a spectroscopic deterministic solution in DR3, although the proportion of systems lost at every filtering step may markedly vary. As an illustration, only the very brightest ASAS-SN eclipsing binaries get a DR3 solution. While the range is not an issue, we find that 90% have a (combined) magnitude fainter than 12 mag. Similar figures apply to the WISE sample. Regardless of these intrinsic limitations, it is naturally expected that the completeness level will substantially improve in DR4 thanks to the larger amount of RVs collected and improvements in data analysis.
@@ 10.1.1 Intrinsic variables
Since the NSS-SB1 processing is applied to time-series of variable RVs, it is expected to also treat some intrinsic variables that are exhibiting well-defined RV variations. This could be the case for example for Cepheids (actually the bright ones). The only firewall preventing the processing to deliver spurious orbital solutions for intrinsic variables is that the processing only considers an harmonic content of the RV curve that is Keplerian (or nearly so). Unfortunately, Cepheids RV curves, for example, are renowned for their harmonic content being quite similar to the Keplerian one. This has been pointed out by Imbert et al. (1985); Imbert (1987); Imbert et al. (1989). The problem generated by intrinsic variables has been pointed out in Sect. 9.4.4 and in Sect. 9.4.5 where some of them were detected in the binary catalogue, mainly at short periods and high eccentricities.
As part of DR3, Ripepi et al. (2023) published a complete analysis of the Gaia photometric curves for a little more than 15,000 Cepheids (of all classes), 799 having time-series of RVs. This list was cross-correlated with our SB1/SB1C list. They found 338 objects in common. Among them, the Cepheid nature is confirmed from the literature for 276 objects. After an eye inspection of photometric and RV curves, five of them were found doubtful Cepheids and have been rejected. Finally, 333 spurious SB1 solutions due to the Cepheid nature of the objects have been blacklisted from the SB-subcatalogue.
Along these lines, we also considered the case of RR Lyrae stars (of all classes too). Clementini et al. (2023) gathered a list of a little more than 270,000 such objects, among which some 1096 have time-series of RVs. The cross-match with our SB1 list yields 152 objects in common. Among these, 5 turned out to have a discrepancy between the period from our SB1 pipeline and the one derived by Clementini et al. (2023). Their RR Lyrae classification were thus considered as insecure. Therefore, we rejected (blacklisted), from the SB1 list, 147 objects considered as bona fide RR Lyrae. In summary, we rejected from our SB1 list some 480 sources as being intrinsic variables. Although a few pulsating stars (Cepheids, RR Lyrae) are rejected from our SB-subcatalogue, they are nevertheless studied as pulsators elsewhere (see e.g. Eyer et al. 2023).
@@
We will illustrate here our concerns by further studying the case of Gaia DR3 5829160851462523008. This object was not known to vary before Gaia. It appears in the catalogue of LPV candidates with a derived photometric period of 90.5 d. This object also appears in the SB-subcatalogue, produced in the present work, as a binary with a period of 178.886 d. Figure 47 exhibits both the RV curve and the photometric light curve of this object with the phase computed according to the RV period. From this figure, it is clear that the RV curve is unimodal, as it should be, whereas the light curve presents two cycles in good agreement with the period attributed by Lebzelter et al. (2023). If the two maxima are very similar, the two minima are significantly different. In addition, the minima happen at phases where the RV curve crosses the systemic velocity. All this strongly suggests that this star is an ellipsoidal variable. Therefore, at the time of DR3, it appeared hazardous to filter our solutions on the basis of the existing LPV catalogue. Objects like Gaia DR3 5829160851462523008 would have been unduly rejected and we cautiously decided to not reject LPVs on this basis. Certainly, a sound analysis necessitates the concomitant study of RV and photometry, which was not foreseen for DR3. This should be done in the subsequent FPR (see Gaia Collaboration et al. 2023c) where some 20 % of the LPV candidates are expected to be ellipsoidal variables. We encourage the user of the SB-subcatalogue to filter out true intrinsic variables thanks to these subsequent results.
@@
The problem of the rejection of intrinsic variables is certainly a difficult one and much progress is still to be made on this topic. As mentioned above, the measurement of RVs by cross-correlating the spectra with template prevents to identify the origin of the RV variations.
@@ 10.1.3 Scan-angle dependent signals
The RVS being a slitless spectrometer, the knowledge of the exact position of a star is necessary to have a correct calibration of the wavelength for its observed spectrum (Sartoretti et al. 2018). The astrometric solution assumes that all the objects are point-like, and the objects presenting incompatibility with this hypothesis have potentially wrong RVs. They are thus associated to large values of the ruwe parameter. This point was already mentioned in Sect. 2. To these deviating objects is associated an excess noise compared to the global astrometric solution (labelled astrometric_excess_noise). One of the main categories of the non-point-like objects are close pairs (physical or not) unresolved or partially resolved. A blend of objects that presents some elongated image will induce the following effect. The measured position of the star will have a position dependent on how the elongated axis of the pair enters the astrometric CCD field. Thus, the resulting RVs are dependent on the orientation of the main axis of the elongated objects with respect to the scan. It depends on the scan angle in a rather smooth manner. This potentially could generate, from transit to transit, RV variations that are a simple function of the scan angle. In particular, solutions for RV-constant objects could be associated to solutions mimicking a spectroscopic orbit at a period of 62.97 d; these orbits are consequently spurious. During the precession cycle, the scan angles of the various transits are performing a complete round. The phenomenon manifests itself as an excess of solutions with small and thus small mass function (because the period is fixed). It is visible in Fig. 31 and is further illustrated in Fig. 48. Clearly, most of the solutions are spurious and are associated to large values of the astrometric excess noise. The suspicious span the approximate interval between 0.8 and 16 km s-1. Another illustration can be found in Gaia Collaboration et al. (2023a), where it is also shown that the phenomenon could be related to a particular behaviour of ipd_gof_harmonic_amplitude. A detailed investigation of this behaviour is to be found in Holl et al. (2023a).
@@
Thus, the epoch RVs are affected by unmodelled astrometric shifts along the scan direction. In order to alleviate the problem, the objects in the class SB1 were first selected on the basis of an astrometric ruwe larger than 1.4 and a solution leading to a period within 3% of the precession period (62.97 d). In addition, we selected objects obeying the condition
@@
where is the mean of the uncertainties on the individual epoch RVs over the full time-series. Fig. 49 illustrates the distribution of the astrometric_excess_noise as a function of the parameter for the selected sample. One clearly sees that all the values are lower than three times the astrometric_excess_noise suggesting that these orbital solutions are spurious and are due to a scan-angle effect. The semi-amplitude is too small relatively to the error induced by the astrometric solution. It has been decided to reject and blacklist the 87 objects presenting these characteristics.

### v2

@@ 9.2.2 FirstDegreeTrendSB1 and SecondDegreeTrendSB1 classes
Figure 33 illustrates the distribution of the first derivative normalised by its uncertainty. The distribution is very similar for the two classes and presents an almost perfect symmetry, as expected. The deficit of solutions around the zero derivative is due to the fact that these solutions are not included because they correspond to constant objects. A lot of the solutions are well above . A few well-defined derivatives are further away up to .
@@ 9.3 Validation against CU7 photometric periods
Some intrinsic variables can exhibit RV variations that could be mistaken as the effect of an orbital motion. As explained in Sect. 10.1.1, a sample of securely identified Cepheids (333, Ripepi et al. 2023) and RR Lyrae (147, Clementini et al. 2023) have also an SB1 solution. They should be eventually excluded from our SB subcatalogue but they also offer the unique opportunity to have an independent test of the spectroscopic periods by comparing them to the photometric ones. Indeed, we compared our spectroscopic periods to those determined from DR3 photometric data for the 147 RR Lyrae and the 333 Cepheids. The difference is expressed as [P(spectroscopy)–P(photometry)]/, where is the square root of the sum in quadrature of both uncertainties. We find a recovery success rate to within 3 of 95% for the RR Lyrae’s. It is not only lower for the Cepheids (85%), but the outliers are also much more severely discrepant (Fig. 35). We find that the solution efficiency is better for the RR Lyrae’s than for the Cepheids.
@@ 9.4.1 Comparison against Griffin’s results
SB1: 83 systems,
SB1C: 0 system,
TrendSB1: 10 systems,
SB2: 1 system,
SB2C: 0 system,
@@ 9.4.2 Comparison against SB9 catalogue
As a next step, the analysis above was repeated for the whole SB9. Although the sample is now considerably larger, it is much less homogeneous and contains solutions of generally lower quality. The most recent orbit was selected when a given system had multiple entries in the catalogue. As described above, visual and/or multiple systems were also excluded. Six SB2 orbits were replaced by those determined by Lester (2020). This validation catalogue contains the orbits of 1750 SB1’s and 647 SB2’s. The breakdown of DR3 spectroscopic solution types is as follows for SB1s in the SB9 catalogue:
SB1: 226 systems,
SB1C: 0 system,
TrendSB1: 45 systems,
SB2: 3 systems,
SB2C: 0 system,
@@
All (but one) of the 45 systems with a trend solution have very long periods. Furthermore, 18 and 2 systems in the SB2 validation catalogue have received a SB1 and a TrendSB1 solution, respectively.
@@ 9.4.3 Comparison against APOGEE
Orbital parameters for 20 000 SB1’s have been released as part of APO Galactic Evolution Experiment (APOGEE)141414http://www.sdss3.org/surveys/apogee.php DR16 (Price-Whelan et al. 2020, hereafter PW20). Our validation sample is their ’Gold Sample’ that contains 1032 systems with high-quality, unimodal posterior samplings.
The breakdown of DR3 spectroscopic solution types is as follows for the SB1s in PW20:
SB1: 126 systems,
SB1C: 0 system,
TrendSB1: 4 systems,
SB2: 3 systems,
SB2C: 0 system,
@@ 9.4.4 Comparison against WISE
A total of 56 000 short-period ( 10 d) binary systems observed by the Wide-field Infrared Survey Explorer (WISE)151515https://www.jpl.nasa.gov/missions/wide-field-infrared-survey-explorer-wise have been discussed by Petrosky et al. (2021). Most of them are (near-)contact EBs. However, as will be shown below, a non-negligible amount are classical pulsators, especially RR Lyrae stars. The infrared photometric measurements are made through the W1 filter centred at about 3.4 m. A subset of 49 465 systems can be unambiguously associated to a DR3 source161616http://cdn.gea.esac.esa.int/Gaia/gedr3/cross_match/. This sample allows us to test the reliability of our orbital periods for very close binaries, although other orbital parameters are not available.
The breakdown of DR3 spectroscopic solution types is as follows for the SB1s in Petrosky et al. (2021):
SB1: 704 systems,
SB1C: 3 systems,
TrendSB1: 12 systems,
SB2: 69 systems,
SB2C: 23 systems,
Without any DR3 spectroscopic deterministic solutions: 48 654 systems.
@@ 9.4.5 Comparison against ASAS-SN
The physical parameters of 35 464 detached eclipsing binaries (DEBs) observed in the framework of the All-Sky Automated Survey for Supernovae (ASAS-SN; Kochanek et al. 2017)171717https://www.astronomy.ohio-state.edu/asassn/ have been determined by Rowan et al. (2022). A total of 35 307 systems have a Gaia DR3 ID. The data are based on the analysis of - and -band light curves. The orbital periods nicely bridge the gap between the WISE sample of close binaries and other more general catalogues, such as APOGEE or the SB9 (Fig. 36). Contrary to WISE, other orbital elements (e.g. eccentricity) are provided besides the period, although without any uncertainties. For , we assume a typical relative uncertainty of 10-6 (e.g. Holanda et al. 2018).
@@
SB1: 626 systems,
SB1C: 3 systems,
TrendSB1: 6 systems,
SB2: 92 systems,
SB2C: 35 systems,
Without any DR3 spectroscopic deterministic solutions: 34 545 systems.
@@ 9.4.6 Summary
As discussed above, the success rate for the main orbital parameters (, , , and ) is usually well above 80% for the whole set of validation catalogues considered here (Griffin, SB9, APOGEE, WISE, and ASAS-SN). As expected, the performance increases for relatively short-period systems for which at least one orbital period is covered by the observations (about 2.5 years for the RVS spectroscopic processing data set considered here). In this case, the recovery rate commonly exceeds 85-90%. The WISE sample of (near-)contact binaries allows us to investigate the performance of our period search at high frequencies. We recover the WISE photometric periods at a level of about 80-85%. It increases to about 90% for the DEBs with wider orbits discussed by Rowan et al. (2022). We can also note the excellent success rate for the pulsation periods of the RR Lyrae stars that are below one day (Sect. 9.3). The good agreement on the period with both EB data sets advocates for a deeper study using these data. Their combination will certainly result in an extensive analysis that could reveal various kinds of objects such as EW binaries and/or ellipsoidal variables. Such a task is beyond the scope of the present paper. Our analysis globally suggests that the orbital periods below about 10 days are generally robust provided the solution significance is not too low. We find a roughly constant success rate at the level of about 90-95% once the significance exceeds 15-20. However, although they can often be regarded as compatible with zero within the uncertainties, it appears that the eccentricities may be overestimated even if the significance is larger.
@@
In this respect, the fact that only a relatively small fraction of the SB1s in the validation catalogues have a spectroscopic deterministic DR3 solution is the consequence of the various filtering steps, both at CU4 and CU6 levels, described previously. In addition, we note that for the SB9 and Griffin’s sample, only a subset of the reference catalogue was used. The case of the SB9, and the steps that led to the relatively small overlap with respect to DR3 (see Sect. 9.4.2) are discussed in detail in Appendix G. Similar arguments apply to the other external catalogues and explain the limited number of benchmark binaries with a spectroscopic deterministic solution in DR3, although the proportion of systems lost at every filtering step may markedly vary. As an illustration, only the very brightest ASAS-SN eclipsing binaries get a DR3 solution. While the range is not an issue, we find that 90% have a (combined) magnitude fainter than 12 mag. Similar figures apply to the WISE sample. Regardless of these intrinsic limitations, it is naturally expected that the completeness level will substantially improve in DR4 thanks to the larger amount of RVs collected and improvements in data analysis.
@@ 10.1.1 Intrinsic variables
Since the NSS-SB1 processing is applied to time series of variable RVs, it is expected to also treat some intrinsic variables that are exhibiting well-defined RV variations. This could be the case for example for Cepheids (actually the bright ones). The only firewall preventing the processing from delivering spurious orbital solutions for intrinsic variables is the fact that the processing only considers an harmonic content of the RV curve that is Keplerian (or nearly so). Unfortunately, Cepheids RV curves, for example, are renowned for their harmonic content being quite similar to the Keplerian one. This has been pointed out by Imbert et al. (1985); Imbert (1987); Imbert et al. (1989). The problem generated by intrinsic variables has been pointed out in Sect. 9.4.4 and in Sect. 9.4.5 where some of them were detected in the binary catalogue, mainly at short periods and high eccentricities.
As part of DR3, Ripepi et al. (2023) published a complete analysis of the Gaia photometric curves for a little more than 15 000 Cepheids (of all classes), 799 having time series of RVs. This list was cross-correlated with our SB1/SB1C list. They found 338 objects in common. Among them, the Cepheid nature is confirmed from the literature for 276 objects. After an eye inspection of photometric and RV curves, five of them were found doubtful Cepheids and have been rejected. Finally, 333 spurious SB1 solutions due to the Cepheid nature of the objects have been blacklisted from the SB subcatalogue.
Along these lines, we also considered the case of RR Lyrae stars (of all classes too). Clementini et al. (2023) gathered a list of a little more than 270 000 such objects, among which some 1096 have time series of RVs. The cross-match with our SB1 list yields 152 objects in common. Among these, 5 turned out to have a discrepancy between the period from our SB1 pipeline and the one derived by Clementini et al. (2023). Their RR Lyrae classification were thus considered as insecure. Therefore, we rejected (blacklisted), from the SB1 list, 147 objects considered as bona fide RR Lyrae. In summary, we rejected from our SB1 list some 480 sources as being intrinsic variables. Although a few pulsating stars (Cepheids, RR Lyrae) are rejected from our SB subcatalogue, they are nevertheless studied as pulsators elsewhere (see e.g. Eyer et al. 2023).
@@
We illustrate here our concerns by further studying the case of Gaia DR3 5829160851462523008. This object was not known to vary before Gaia. It appears in the catalogue of LPV candidates with a derived photometric period of 90.5 d. This object also appears in the SB subcatalogue, produced in the present work, as a binary with a period of 178.886 d. Figure 47 exhibits both the RV curve and the photometric light curve of this object with the phase computed according to the RV period. From this figure, it is clear that the RV curve is unimodal, as it should be, whereas the light curve presents two cycles in good agreement with the period attributed by Lebzelter et al. (2023). If the two maxima are very similar, the two minima are significantly different. In addition, the minima happen at phases where the RV curve crosses the systemic velocity. All this strongly suggests that this star is an ellipsoidal variable. Therefore, at the time of DR3, it appeared hazardous to filter our solutions on the basis of the existing LPV catalogue. Objects like Gaia DR3 5829160851462523008 would have been unduly rejected and we cautiously decided to not reject LPVs on this basis. Certainly, a sound analysis necessitates the concomitant study of RV and photometry, which was not foreseen for DR3. This should be done in the subsequent FPR (see Gaia Collaboration et al. 2023c) where some 20 % of the LPV candidates are expected to be ellipsoidal variables. We encourage the user of the SB subcatalogue to filter out true intrinsic variables thanks to these subsequent results.
@@
The problem of the rejection of intrinsic variables is certainly a difficult one, and much progress is still to be made on this topic. As mentioned above, the measurement of RVs by cross-correlating the spectra with template prevents us from identifying the origin of the RV variations.
@@ 10.1.3 Scan-angle dependent signals
The RVS being a slitless spectrometer, the knowledge of the exact position of a star is necessary to have a correct calibration of the wavelength for its observed spectrum (Sartoretti et al. 2018). The astrometric solution assumes that all the objects are point-like, and the objects presenting incompatibility with this hypothesis have potentially wrong RVs. They are thus associated to large values of the ruwe parameter. This point is also mentioned in Sect. 2. To these deviating objects is associated an excess noise compared to the global astrometric solution (labelled astrometric_excess_noise). One of the main categories of the non-point-like objects are close pairs (physical or not) unresolved or partially resolved. A blend of objects that presents some elongated image will induce the following effect. The measured position of the star will have a position dependent on how the elongated axis of the pair enters the astrometric CCD field. Thus, the resulting RVs are dependent on the orientation of the main axis of the elongated objects with respect to the scan. It depends on the scan angle in a rather smooth manner. This potentially could generate, from transit to transit, RV variations that are a simple function of the scan angle. In particular, solutions for RV-constant objects could be associated to solutions mimicking a spectroscopic orbit at a period of 62.97 d; these orbits are consequently spurious. During the precession cycle, the scan angles of the various transits are performing a complete round. The phenomenon manifests itself as an excess of solutions with small and thus small mass function (because the period is fixed). It is visible in Fig. 31 and is further illustrated in Fig. 48. Clearly, most of the solutions are spurious and are associated to large values of the astrometric excess noise. The suspicious span the approximate interval between 0.8 and 16 km s-1. Another illustration can be found in Gaia Collaboration et al. (2023a), where it is also shown that the phenomenon could be related to a particular behaviour of ipd_gof_harmonic_amplitude. A detailed investigation of this behaviour is to be found in Holl et al. (2023a).
@@
Thus, the epoch RVs are affected by unmodelled astrometric shifts along the scan direction. In order to alleviate the problem, the objects in the class SB1 were first selected on the basis of an astrometric ruwe larger than 1.4 and a solution leading to a period within 3% of the precession period (62.97 d). In addition, we selected objects obeying the condition:
@@
where is the mean of the uncertainties on the individual epoch RVs over the full time series. Figure 49 illustrates the distribution of the astrometric_excess_noise as a function of the parameter for the selected sample. We can clearly see that all the values are lower than three times the astrometric_excess_noise suggesting that these orbital solutions are spurious and are due to a scan-angle effect. The semi-amplitude is too small relatively to the error induced by the astrometric solution. It has been decided to reject and blacklist the 87 objects presenting these characteristics.